# Exactly solvable Schrödinger operators related to the hypergeometric equation

### v1

@@ Abstract
We study one-dimensional Schrödinger operators defined as closed operators that are exactly solvable in terms of the Gauss hypergeometric function. We allow the potentials to be complex.
These operators fall into three groups. The first group can be reduced to the Gegenbauer equation, up to an affine transformation, a special case of the hypergeometric equation. The two other groups, which we call hypergeometric of the first, resp. second kind, can be reduced to the general Gauss hypergeometric equation. Each of the group is subdivided in three families, acting to on the Hilbert space $L^{2}]-1,1[,$ $L^{2}({\mathbb{R}}_{+})$ resp. $L^{2}({\mathbb{R}})$. Motivated by geometric applications of these families, we call them spherical, hyperbolic, resp. deSitterian. All these families are known from applications in Quantum Mechanics: e.g. spherical hypergeometric Schrödinger operators of the first kind are often called trigonometric Pöschl-Teller Hamiltonians.
For operators belonging to each family we compute their spectrum and determine their Green function (the integral kernel of their resolvent). We also describe transmutation identities that relate these Green functions. These identities interchange spectral parameters with coupling constants across different operator families. Finally, we describe how these operators arise from separation of variables of (pseudo-)Laplacians on symmetric manifolds.
Our paper can be viewed as a sequel to [DL], where closed realizations of one-dimensional Schrödinger operators solvable in terms Kummer’s confluent equation were studied.
@@ 1.2 Review of 9 families
      $L_{\alpha}^{\rm s}$ $:=-\partial_{r}^{2}+\left(\alpha^{2}-\frac{1}{4}\right)\frac{1}{\sin^{2}r},$ (1.4)
@@
      $L_{\alpha}^{\rm h}$ $:=-\partial_{r}^{2}+\left(\alpha^{2}-\frac{1}{4}\right)\frac{1}{\sinh^{2}r},$ (1.5)
@@
      $L_{\alpha}^{\mathrm{dS}}$ $:=-\partial_{r}^{2}-\left(\alpha^{2}-\frac{1}{4}\right)\frac{1}{\cosh^{2}r},$ (1.6)
@@
      $L_{\alpha,\beta}^{\rm s}$ $:=-\partial_{r}^{2}+\left(\alpha^{2}-\frac{1}{4}\right)\frac{1}{4\sin^{2}\frac{r}{2}}+\left(\beta^{2}-\frac{1}{4}\right)\frac{1}{4\cos^{2}\frac{r}{2}},$ (1.7)
@@
      $L_{\alpha,\beta}^{\rm h}$ $:=-\partial_{r}^{2}+\left(\alpha^{2}-\frac{1}{4}\right)\frac{1}{4\sinh^{2}\frac{r}{2}}-\left(\beta^{2}-\frac{1}{4}\right)\frac{1}{4\cosh^{2}\frac{r}{2}},$ (1.8)
@@
      $L_{\alpha,\beta}^{\mathrm{dS}}$ $:=-\partial_{r}^{2}-\left(\alpha^{2}-\frac{1}{4}\right)\frac{1}{\cosh^{2}r}\left(\frac{1}{2}+\frac{{\rm i}\sinh r}{2}\right){-}\left(\beta^{2}-\frac{1}{4}\right)\frac{1}{\cosh^{2}r}\left(\frac{1}{2}-\frac{{\rm i}\sinh r}{2}\right),$ (1.9)
@@
      $K_{\tau,\mu}^{\rm s}$ $:=-\partial_{u}^{2}+\tau\frac{\cos u}{\sin u}+\left(\frac{\mu^{2}}{4}-\frac{1}{4}\right)\frac{1}{\sin^{2}u},$ (1.10)
@@
      $K_{\kappa,\mu}^{\rm h}$ $:=-\partial_{u}^{2}+\kappa\frac{\cosh u}{\sinh u}+\left(\frac{\mu^{2}}{4}-\frac{1}{4}\right)\frac{1}{\sinh^{2}u},$ (1.11)
@@
      $K_{\kappa,\mu}^{\mathrm{dS}}$ $:=-\partial_{u}^{2}-\kappa\frac{\sinh w}{\cosh w}-\left(\frac{\mu^{2}}{4}-\frac{1}{4}\right)\frac{1}{\cosh^{2}w},$ (1.12)
@@
Gegenbauer Hamiltonians are special cases of both hypergeometric Hamiltonians of the first and second type. In fact, we have the following coincidencies:
@@ 1.4 Geometric applications
All three types of hypergeometric Hamiltonians of the 1st kind have natural geometric interpretations as well. In the following list we show how they arise from separtion of variables in a (pseudo-)Laplacian on a (pseudo-)sphere in “double spherical coordinates”. In all three examples the coordinates in the ambient pseudo-Euclidean space are partitioned into two groups.
@@
    $\Big{(}\sin\frac{r}{2}\Big{)}^{\frac{p-1}{2}}\Big{(}\cos\frac{r}{2}\Big{)}^{\frac{q-1}{2}}\big{(}-\Delta_{p+q-1}^{\rm s}\big{)}\Big{(}\sin\frac{r}{2}\Big{)}^{-\frac{p-1}{2}}\Big{(}\cos\frac{r}{2}\Big{)}^{-\frac{q-1}{2}}+\big{(}\tfrac{p+q-2}{2}\big{)}^{2}$ (1.46)
    $=$ $4\Bigg{(}-\partial_{r}^{2}+\frac{\big{(}\tfrac{p-2}{2}\big{)}^{2}-\tfrac{1}{4}-\Delta_{p-1}^{\rm s}}{4\sin^{2}\frac{r}{2}}+\frac{\big{(}\tfrac{q-2}{2}\big{)}^{2}-\tfrac{1}{4}-\Delta_{q-1}^{\rm s}}{4\cos^{2}\frac{r}{2}}\Bigg{)}.$ (1.47)
@@
    $\Big{(}\cosh\frac{r}{2}\Big{)}^{\frac{p-1}{2}}\Big{(}\sinh\frac{r}{2}\Big{)}^{\frac{q-1}{2}}\big{(}-\Delta_{p-1,q}\big{)}\Big{(}\cosh\frac{r}{2}\Big{)}^{-\frac{p-1}{2}}\Big{(}\sinh\frac{r}{2}\Big{)}^{-\frac{q-1}{2}}-\big{(}\tfrac{p+q-2}{2}\big{)}^{2}$ (1.48)
    $=$ $4\Bigg{(}-\partial_{r}^{2}-\frac{\big{(}\tfrac{p-2}{2}\big{)}^{2}-\tfrac{1}{4}-\Delta_{p-1}^{\rm s}}{4\cosh^{2}\frac{r}{2}}+\frac{\big{(}\tfrac{q-2}{2}\big{)}^{2}-\tfrac{1}{4}-\Delta_{q-1}^{\rm s}}{4\sinh^{2}\frac{r}{2}}\Bigg{)}.$ (1.49)
@@
    $\big{(}\cosh r)^{\frac{p-1}{2}}\big{(}\Delta_{p-1,p}\big{)}(\cosh r)^{-\frac{p-1}{2}}-(p-1)^{2}$ (1.50)
@@
We use the geometric interpretation of Gegenbauer Hamiltonians as the main justification for our names of types: spherical, hyperbolic and deSitterian. For coherence, we extend these names to hypergeometric Hamiltonians of both kinds. The names “spherical” and “hyperbolic” seem quite non-controversial, and are used in the literature in similar contexts. The name “deSitterian” is our invention. Its justification is somewhat weaker and based only on the Gegenbauer Hamiltonian: the connection of deSitterian hypergeometric Hamiltonians of the first kind and de Sitter spaces is less obvious.
@@ 1.5 Boundary conditions
    Short range potential. $b=+\infty$ and $V(x)$ is integrable near $\infty$. Then $\nu_{b}(L)=0.$ Moreover, eigenfunctions in $\mathcal{D}(L_{\bullet})$ with eigenvalue $k^{2}$, ${\rm Re}(k)>0$ behave as ${\rm e}^{-kx}$ with ${\rm Re}(k)>0$.
@@
    Shifted short range potential. $b=+\infty$ and $V(x)$ is a constant plus integrable near $\infty$. Same as before, except that the eigenfunctions $\sim{\rm e}^{-kx}$ have energy $k^{2}+V(+\infty)$.
  3. 3.
@@
    Then $\nu_{b}(L)=2$ iff $|{\rm Re}(m)|<1$, otherwise $\nu_{b}(L)=0$. The behavior of elements of $\mathcal{D}(L_{\bullet})$ near $b$ is $|x-b|^{\frac{1}{2}+m}$ for ${\rm Re}(m)\geq 1$, a linear combination of $|x-b|^{\frac{1}{2}+m}$ and $|x-b|^{\frac{1}{2}-m}$ for $0\leq{\rm Re}\,m<1$, $m\neq 0$ and a linear combination of $|x-b|^{\frac{1}{2}}$ and $|x-b|^{\frac{1}{2}}\ln|x-b|$ for $m=0$. Note that $m=\frac{1}{2}$ corresponds to the Dirichlet b.c. and $m=-\frac{1}{2}$ to the Neumann b.c.
@@
  4. 4.
@@
Spherical of the first kind Bessel Bessel
Hyperbolic of the first kind Bessel Short range potential
Desitterian of the first kind Short range potential Short range potential
Spherical of the second type Whittaker Whittaker
Hyperbolic of the second type Whittaker Short range potential
Desitterian of the second type Shifted short range potential Shifted short range potential
@@ 1.6 Comparison with literature and historic remarks
which acts on the Hilbert space $L^{2}\big{(}]-1,1[,(1-x)^{\alpha}(1+x)^{\beta})$ see e.g. [GPLS] and [Koo]. Its eigenfunctions are the famous Jacobi polynomials. By a simple transformation the Jacobi operator is unitarily equivalent to trigonometric Pöschl-Teller Hamiltonian (which we call spherical hypergeometric Hamiltonian of the first kind).
@@
An interesting review of various exactly solvable Schrödinger operators is contained in [Everitt]. It does not, however, contain all 9 families that we consider.
@@
Let us briefly outline the history of these Hamiltonians in physics here. In the physics community, the study of Schrödinger operators exactly solvable in terms of hypergeometric functions started in the 1930’s when physicists studied biatomic and polyatomic molecules dynamics and the exact solution to the Schrödinger equation. The main reason these physicists considered such Hamiltonians was often due to the limitations of the perturbation method. Therefore, various kinds of exactly solvable Hamiltonians were suggested which fit the experimental data. That was the primary motivation behind the Hamiltonians proposed by Morse [Mor], Rosen-Morse [MoR], Eckart [Eck], and Manning-Rosen [MaR].
@@
The geometric interpretation of hypergeometric Hamiltonians is closely related to the analysis of hypergeometric equation based on Lie groups and Lie algebras, which possess large literature [M1, V, Wa, D2]. The interpretation of the deSitterian hypergeometric Hamiltonian of the 1st kind in terms of complex spheres in 1.50 seems to be new.
@@ 1.7 Plan of the paper
The core of the paper are the sections 3, 4 and 5, where the $3\times 3$ families of Hamiltonians are introduced and studied. For each member of each family its spectrum and Green functions are computed. We also prove various identities that we already listed in the introduction, including the transmutation identities.
@@
In Appendix A we collect identities about hypergeometric and Gegenbauer functions that we use in our paper.
In Appendix B we give a very concise account of the theory of closed realizations of 1d Schrödinger operators. This account is incomplete and not fully rigorous—the reader should consult Sect. 2 of [DL] for a more detailed and rigorous exposition, or [DeGe], where a complete theory with proofs is given. Of course, the topic is classic and contained in other texts such as [DuSch, EE, DeGe].
In our paper we try to use notation and conventions that make our formulas, especially for Green functions, as simple, elegant and symmetric as possible. This often motivates us to introduce our conventions, different from the standard ones. In particular, we do not use the standard conventions for associated Legendre functions, which could be used for Gegenbauer Hamiltonians. For readers used to associated Legendre functions, in Appendix C we recall their definitions and compare them with the special functions that we use.

### v2

@@ Abstract
We study one-dimensional Schrödinger operators defined as closed operators that are exactly solvable in terms of the Gauss hypergeometric function. We allow the potentials to be complex. These operators fall into three groups. The first group can be reduced to the Gegenbauer equation, up to an affine transformation, a special case of the hypergeometric equation. The two other groups, which we call hypergeometric of the first, resp. second kind, can be reduced to the general Gauss hypergeometric equation. Each of the group is subdivided in three families, acting on the Hilbert space $L^{2}]-1,1[,$ $L^{2}({\mathbb{R}}_{+})$ resp. $L^{2}({\mathbb{R}})$. Motivated by geometric applications of these families, we call them spherical, hyperbolic, resp. deSitterian. All these families are known from applications in Quantum Mechanics: e.g. spherical hypergeometric Schrödinger operators of the first kind are often called trigonometric Pöschl-Teller Hamiltonians. For operators belonging to each family we compute their spectrum and determine their Green function (the integral kernel of their resolvent). We also describe transmutation identities that relate these Green functions. These identities interchange spectral parameters with coupling constants across different operator families. Finally, we describe how these operators arise from separation of variables of (pseudo-)Laplacians on symmetric manifolds. Our paper can be viewed as a sequel to [DL], where closed realizations of one-dimensional Schrödinger operators solvable in terms Kummer’s confluent equation were studied.
@@ 1.2 Review of 9 families
Note that we always use the notion of an analytic family of closed operators in the sense of Kato [Kato, DW1]. In Appendix A we give a concise explanation of this concept.
@@
      $L_{\alpha}^{\rm s}$ $=-\partial_{r}^{2}+\left(\alpha^{2}-\frac{1}{4}\right)\frac{1}{\sin^{2}r},$ (1.4)
@@
      $L_{\alpha}^{\rm h}$ $=-\partial_{r}^{2}+\left(\alpha^{2}-\frac{1}{4}\right)\frac{1}{\sinh^{2}r},$ (1.5)
@@
      $L_{\alpha}^{\mathrm{dS}}$ $=-\partial_{r}^{2}-\left(\alpha^{2}-\frac{1}{4}\right)\frac{1}{\cosh^{2}r},$ (1.6)
@@
      $L_{\alpha,\beta}^{\rm s}$ $=-\partial_{r}^{2}+\left(\alpha^{2}-\frac{1}{4}\right)\frac{1}{4\sin^{2}\frac{r}{2}}+\left(\beta^{2}-\frac{1}{4}\right)\frac{1}{4\cos^{2}\frac{r}{2}},$ (1.7)
@@
      $L_{\alpha,\beta}^{\rm h}$ $=-\partial_{r}^{2}+\left(\alpha^{2}-\frac{1}{4}\right)\frac{1}{4\sinh^{2}\frac{r}{2}}-\left(\beta^{2}-\frac{1}{4}\right)\frac{1}{4\cosh^{2}\frac{r}{2}},$ (1.8)
@@
      $L_{\alpha,\beta}^{\mathrm{dS}}$ $=-\partial_{r}^{2}-\left(\alpha^{2}-\frac{1}{4}\right)\frac{1}{\cosh^{2}r}\left(\frac{1}{2}+\frac{{\rm i}\sinh r}{2}\right){-}\left(\beta^{2}-\frac{1}{4}\right)\frac{1}{\cosh^{2}r}\left(\frac{1}{2}-\frac{{\rm i}\sinh r}{2}\right),$ (1.9)
@@
      $K_{\tau,\mu}^{\rm s}$ $=-\partial_{u}^{2}+\tau\frac{\cos u}{\sin u}+\left(\frac{\mu^{2}}{4}-\frac{1}{4}\right)\frac{1}{\sin^{2}u},$ (1.10)
@@
      $K_{\kappa,\mu}^{\rm h}$ $=-\partial_{u}^{2}+\kappa\frac{\cosh u}{\sinh u}+\left(\frac{\mu^{2}}{4}-\frac{1}{4}\right)\frac{1}{\sinh^{2}u},$ (1.11)
@@
      $K_{\kappa,\mu}^{\mathrm{dS}}$ $=-\partial_{u}^{2}-\kappa\frac{\sinh w}{\cosh w}-\left(\frac{\mu^{2}}{4}-\frac{1}{4}\right)\frac{1}{\cosh^{2}w},$ (1.12)
@@
Gegenbauer Hamiltonians are special cases of hypergeometric Hamiltonians of both the first and second type. In fact, we have the following coincidencies:
@@ 1.3 Transmutations of Green functions
In the literature the name “transmutation” is sometimes used for identities of the form $A=VBV^{-1}$, e.g. in [Car, SS]. Note that our transmutation identities are not of this form, and in fact they seem to represent a novel contribution.
@@ 1.4 Geometric applications
All three types of hypergeometric Hamiltonians of the 1st kind have natural geometric interpretations as well. In the following list we show how they arise from separation of variables in a (pseudo-)Laplacian on a (pseudo-)sphere in “double spherical coordinates”. In all three examples the coordinates in the ambient pseudo-Euclidean space are partitioned into two groups.
@@
    $\left(\sin\frac{r}{2}\right)^{\frac{p-1}{2}}\left(\cos\frac{r}{2}\right)^{\frac{q-1}{2}}\left(-\Delta_{p+q-1}^{\rm s}\right)\left(\sin\frac{r}{2}\right)^{-\frac{p-1}{2}}\left(\cos\frac{r}{2}\right)^{-\frac{q-1}{2}}+\left(\tfrac{p+q-2}{2}\right)^{2}$
    $=$ $4\left(-\partial_{r}^{2}+\frac{\left(\tfrac{p-2}{2}\right)^{2}-\tfrac{1}{4}-\Delta_{p-1}^{\rm s}}{4\sin^{2}\frac{r}{2}}+\frac{\left(\tfrac{q-2}{2}\right)^{2}-\tfrac{1}{4}-\Delta_{q-1}^{\rm s}}{4\cos^{2}\frac{r}{2}}\right).$ (1.46)
@@
    $\left(\cosh\frac{r}{2}\right)^{\frac{p-1}{2}}\left(\sinh\frac{r}{2}\right)^{\frac{q-1}{2}}\left(-\Delta_{p-1,q}\right)\left(\cosh\frac{r}{2}\right)^{-\frac{p-1}{2}}\left(\sinh\frac{r}{2}\right)^{-\frac{q-1}{2}}-\left(\tfrac{p+q-2}{2}\right)^{2}$
    $=$ $4\left(-\partial_{r}^{2}-\frac{\left(\tfrac{p-2}{2}\right)^{2}-\tfrac{1}{4}-\Delta_{p-1}^{\rm s}}{4\cosh^{2}\frac{r}{2}}+\frac{\left(\tfrac{q-2}{2}\right)^{2}-\tfrac{1}{4}-\Delta_{q-1}^{\rm s}}{4\sinh^{2}\frac{r}{2}}\right).$ (1.47)
@@
    $\left(\cosh r\right)^{\frac{p-1}{2}}\left(\Delta_{p-1,p}\right)\left(\cosh r\right)^{-\frac{p-1}{2}}-(p-1)^{2}$
    $=$ $4\left(-\partial_{r}^{2}-\frac{\left(-\Delta_{p-1}^{{\rm s},{\mathbb{C}}}+\left(\frac{p-1}{2}\right)^{2}-\frac{1}{4}\right)}{2(1+{\rm i}\sinh r)}-\frac{\left(-\overline{\Delta_{p-1}^{{\rm s},{\mathbb{C}}}}+\left(\frac{p-1}{2}\right)^{2}-\frac{1}{4}\right)}{2(1-{\rm i}\sinh r)}\right).$ (1.48)
@@
We use the geometric interpretation of Gegenbauer Hamiltonians as the main justification for our names of types: spherical, hyperbolic and deSitterian. For coherence, we extend these names to hypergeometric Hamiltonians of both kinds. The names “spherical” and “hyperbolic” seem quite non-controversial, and are used in the literature in similar contexts. The name “deSitterian” is our invention. Its justification is somewhat weaker and based on the geometric interpretation of the Gegenbauer Hamiltonian, see (1.45). The natural geometric interpretation of “deSitterian hypergeometric Hamiltonians of the first kind”, which we described in (1.48) involves the hyperboloid ${\mathbb{H}}^{p-1,p}$, which is a deSitter space only for $p=2$.
@@ 1.5 Boundary conditions
    Short range potential. $b=+\infty$ and $V(x)$ is integrable near $\infty$ or constant plus integrable. Then $\nu_{b}(L)=0.$ Moreover, eigenfunctions in $\mathcal{D}(L_{\bullet})$ with eigenvalue $k^{2}+V(+\infty)$, ${\rm Re}(k)>0$ behave as ${\rm e}^{-kx}$ with ${\rm Re}(k)>0$.
@@
    Then $\nu_{b}(L)=2$ iff $|{\rm Re}\,m|<1$, otherwise $\nu_{b}(L)=0$. The behavior of elements of $\mathcal{D}(L_{\bullet})$ near $b$ is $|x-b|^{\frac{1}{2}+m}$ for ${\rm Re}\,m\geq 1$, a linear combination of $|x-b|^{\frac{1}{2}+m}$ and $|x-b|^{\frac{1}{2}-m}$ for $0\leq{\rm Re}\,m<1$, $m\neq 0$, and a linear combination of $|x-b|^{\frac{1}{2}}$ and $|x-b|^{\frac{1}{2}}\ln|x-b|$ for $m=0$.
@@
    Note that $|x-b|^{\frac{1}{2}+m}$ for $m=\frac{1}{2}$ is |xb||x-b|, and for $m=-\frac{1}{2}$ is $|x-b|^{0}$. Therefore, $m=\frac{1}{2}$ corresponds to the Dirichlet b.c. and $m=-\frac{1}{2}$ to the Neumann b.c. We discuss separately these cases in Sect. 6.
  3. 3.
@@
Spherical of first kind Bessel Bessel
Hyperbolic of first kind Bessel Short range potential
DeSitterian of first kind Short range potential Short range potential
Spherical of second type Whittaker Whittaker
Hyperbolic of second type Whittaker Short range potential
DeSitterian of second type Short range potential Short range potential
@@ 1.6 Comparison with literature and historic remarks
which acts on the Hilbert space $L^{2}\big(]-1,1[,(1-x)^{\alpha}(1+x)^{\beta})$ see e.g. [GPLS, Koo]. Its eigenfunctions are the famous Jacobi polynomials. By a simple transformation the Jacobi operator is unitarily equivalent to the trigonometric Pöschl-Teller Hamiltonian (which we call spherical hypergeometric Hamiltonian of the first kind).
@@
An interesting review of various exactly solvable Schrödinger operators is contained in [Eve]. It does not, however, contain all 9 families that we consider.
@@
Let us briefly outline the history of these Hamiltonians in physics. In the physics community, the study of Schrödinger operators exactly solvable in terms of hypergeometric functions started in the 1930’s when physicists studied biatomic and polyatomic molecules dynamics and the exact solution to the Schrödinger equation. The main reason these physicists considered such Hamiltonians was often due to the limitations of the perturbation method. Therefore, various kinds of exactly solvable Hamiltonians were suggested which fit the experimental data. That was the primary motivation behind the Hamiltonians proposed by Morse [Mor], Rosen-Morse [MoR], Eckart [Eck], and Manning-Rosen [MaR].
@@
Table 2: Names of hypergeometric Hamiltonians appearing in the literature.
Our suggestion Name in [DW] Name in [CKS] Alternative names
Spherical of 1st kind Trigonometric Pöschl–Teller Scarf I Pöschl–Teller of 1st kind, Trigonometric Scarf
Hyperbolic of 1st kind Hyperbolic Pöschl–Teller Scarf II Pöschl–Teller of 2nd kind
DeSitterian of 1st kind Scarf Generalized Pöschl–Teller Hyperbolic Scarf
Spherical of 2nd kind Rosen–Morse Rosen–Morse I Trigonometric Rosen–Morse
Hyperbolic of 2nd kind Eckart Eckart Generalized Morse, Hulthén
DeSitterian of 2nd kind Manning–Rosen Rosen–Morse II Hyperbolic Rosen–Morse, Woods–Saxon
The geometric interpretation of hypergeometric Hamiltonians is closely related to the analysis of hypergeometric equation based on Lie groups and Lie algebras, which possess large literature [M1, V, Wa, D2]. The interpretation of the deSitterian hypergeometric Hamiltonian of the 1st kind in terms of complex spheres in 1.48 seems to be new.
@@ 1.7 Plan of the paper
The core of the paper are the sections 3, 4 and 5, where the $3\times 3=9$ families of Hamiltonians are introduced and studied. For each member of each family its spectrum and Green functions are computed. We also prove various identities that we already listed in the introduction, including the transmutation identities. These results are described in $9$ theorems, which have similar proofs, based mostly on the abstract theory described in Appendix D, and more fully in [DL, DeGe]. The proof of the first theorem, Thm 3.1, is spelled out in more detail, the next ones are more concise.
@@
In Appendix A we recall the definition of an analytic family of closed operators following [Kato, DW1].
In Appendix B we give a concise introduction to differential equations in the complex domain and to the Riemann equation (also called the Papperitz equation). It is very useful to view the Gauss hypergeometric equation as a special case of the Riemann equation. We introduce the notation for the Riemann operator, which is directly inspired by the classic notation due to Papperitz [WW].
In Appendix C we collect identities about hypergeometric and Gegenbauer functions that we use in our paper.
In Appendix D we give a concise account of the theory of closed realizations of 1d Schrödinger operators. This account is incomplete and not fully rigorous—the reader should consult Sect. 2 of [DL] for a more detailed and rigorous exposition, or [DeGe], where a complete theory with proofs is given. Of course, the topic is classic and contained in other texts such as [DuSch, EE, DeGe].
In our paper we try to use notation and conventions that make our formulas, especially for Green functions, as simple, elegant and symmetric as possible. This often motivates us to introduce our conventions, different from the standard ones. In particular, we do not use the standard conventions for associated Legendre functions, which could be used for Gegenbauer Hamiltonians. For readers used to associated Legendre functions, in Appendix E we recall their definitions and compare them with the special functions that we use.